A sulphuric acid has pH=2.Its molarity is a)1/100 b)1/50 c)1/200 d)1/2

Grade:11

1 Answers

pawpawan
6 years ago
ph= -log(H+)2 = -log(H+)Taking anti log on both sides, we get(H+)=1/100As Sulfuric acid is strong acid that is why its dissociation constant is 1.Means no. Moles Of H+ ions = no. Of moles of H2SO4. Hence ans is 1/100